OpenStudy (anonymous):

Braden improves his score with each practice quiz he takes. The number of incorrect answers can be represented by the expression 20 – 5n where n represents the number of practice quizzes he has taken. How many answers will he get incorrect after taking 2 practice quizzes?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{T}}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{n}}\color{blue}{\normalsize\text{k}}\color{purple}{\normalsize\text{s}}\color{purple}{\normalsize\text{!}}\color{red}{\normalsize\text{!}}\color{orange}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

If n = 2, then 20 - 5(2) = 20 - 10 = 10